OpenStudy (anonymous):

A coin was randomly drawn from a bag and then replaced. After 200 draws, it was found that 22 pennies, 53 nickels, 87 dimes and 38 quarters had been drawn. Estimate the probability of drawing a penny.

OpenStudy (crashonce):

11%

OpenStudy (crashonce):

22/200 = 11/100 = 11%

OpenStudy (anonymous):

How did you get that?

OpenStudy (crashonce):

well, u drew 22 out of the 200 so tis just the percent